OpenStudy (anonymous):

3x^2=x±4 How do I solve this?

OpenStudy (phi):

it looks like you have two equations: \[ 3x^2=x+ 4 \\ 3x^2=x- 4 \] solve them individually put the in standard form example \[ 3x^2-x-4=0 \]

OpenStudy (phi):

can you factor that ?